OpenStudy (anonymous):

Solve: 4^(2x-5)<=3^(x-3)

OpenStudy (anonymous):

i'm not sure of what i should do first

OpenStudy (anonymous):

do you mean \[4^{2x-5}=3^{x-3}\] ?

OpenStudy (anonymous):

no. i meant \[\le\] sorry

OpenStudy (anonymous):

Are you doing logarithms or is this just an exponential problem?

OpenStudy (anonymous):

i'm working with logarithms. i just don't know what to do next

OpenStudy (anonymous):

Okay great. I am too. What you have to do is make both sides have the same base. This isn't usually possible but it can be done using log. What you do is rewrite the problem as \[\log 4^{2x-5}=\log 3^{x-5}\] Can you take it from there?

OpenStudy (anonymous):

Replace the equal sign with\[\le\] sorry

OpenStudy (anonymous):

\[2x - 5 \log_44 = x \log_43 - 5\log_43\] Now solve it \(\log_44 = 1\)

OpenStudy (anonymous):

I don't understand this, why is everything in log 4?

OpenStudy (anonymous):

Because I chose to

OpenStudy (anonymous):

It's my choice I could have choose \(\log_3\) as well

OpenStudy (anonymous):

or Natural Log \(\ln\)